How do you know what the period is when graphing sins and cos?
Is the period Y=cos(4X) 2pi?

Answer by ewatrrr(24785)


Hi
cos(x) has a period of 2pi and and amplitude of 1. (Green)
y = acos(bx) has a period of 2pi/b and an amplitude of a
y = cos(4x) has a period of 2pi/4 or pi/2 (Blue)
Note: the Green line extending over the period of 2pi for cos(x)
2*(3.14)=6.28 Green Line extends from (0,1) to (6.28,1)
graph ofcos(x)

graph of cos(4x) Period of pi/2 = 1.57. Blue Line extends from (0,1) to (1.57,1)